OpenStudy (anonymous):

Could someone graph this system of equation? 2x + 3y = –3 x – y = –4

OpenStudy (mathteacher1729):

go to http://www.wolframalpha.com and type plot 2x + 3y = –3 and x – y = –4

OpenStudy (anonymous):

2x + 3y = –3 x – y = –4 let do first one first 1) set x=0 3y=-3 y=-1 (0,-1) set y=0 2x=-3 x= -3/2 (-3/2,0) connect these two point

OpenStudy (mathteacher1729):

Here is how you can graph it without a calc or computer: SUPER IMPORTANT IDEA -- you need only two points to make a line. Let's start with: 2x + 3y = –3 Let x = 0 2(0) + 3y = -3 3y = -3 divide by 3 y = -1 One point on this line is (0,-1) Now let y = 0 2x + 3(0) = -3 2x = -3 divide by 2 x = -3/2 Another point on this line is (-3/2, 0) Now connect the two points (0,-1) and (-3/2, 0) with a ruler. :) Do the same for the other equation. (start with letting x = 0 first and solve for y, then let y = 0 and solve for x.)